Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Trying to understand probabilities better...

Expert replies
by mayonnai5e » Wed Sep 12, 2007 3:24 pm
I can solve basic probability questions and even some more advanced ones, but the problem is it takes too long because I often start off on the wrong track and get confused by how to set up the probability formula. Sometimes I try to multiply several probabilities and other times I try to use the standard formula, but use the wrong calculations. I have a few questions that I hope some of the probability gurus on this forum can help me with and some examples to look at (examples were posted in other threads):
From a group of 3 boys and 3 girls, 4 children are to be randomly selected. What is the probability that equal numbers of boys and girls will be selected

A. 1/10
B. 4/9
C. 1/2
D. 3/5
E. 2/3

If the ans is D ie 3/5 then my explanation goes like this:-
Total no of events = Selecting 4 children from a group of 6
= 6C4 = 15
Favorable events will only be when 2 boys and two girls are selected (Since equal nos)
Selecting 2 Girls from a group of 3 Girls = 3C2 = 3
Selecting 2 Boys from a group of 4 Boys = 3C2 = 3
Total favorable events = 3x3 = 9
Prob = 9/15
or 3/5
In this case, the solution involved using the std prob formula with the combination formula.
Tanya prepared 4 different letters to be sent to 4 different addresses. For each letter, she prepared an envelope with its correct address. If the 4 letters are to be put into the 4 envelopes at random, what is the probability that only 1 letter will be put into the envelope with the correct address?

a) 1/24
b) 1/8
c) 1/4
d) 1/3
e) 3/8

Probability that only 1 letter will be put into the envelope with the correct address is (d) 1/3

Step 1 - No. of ways you can select one letter (out of the four given) = 4
and there is only one way to put this letter into the correct envelope.

Now we need to ensure that none of the other letters (B,C and D) are put into the correct envelopes.

Step 2 - So, the next letter, can be put in an incorrect envelope in 2! ways (because one envelope out of the remaining three envelopes will be the correct one for this letter and hence we cannot use that one)

Step 3 - Now there are two letters remaining and there is only one way to put them into incorrect envelopes.

Required Probability = (4 * 2)/ 4! [Total number of ways is 4!]
= 1/3
In this second case, the std formula is used with permutations.

So my first question is, how can I easily determine whether to use combinations or permutations?

My next question is how not to get off track. I sometimes make this mistake (another forum member posted this in response to the first question:
Ok may be I am just getting very confused but here is what I did and I got the ans as 1/10

To get equal no of boys and girls u need to find the probability of selecting 2 girls and the probability of selecting 2 boys.

Probability of selecting a girl on the first chance is 3/6

and the prob of selecting a girl in the 2nd chance is 2/5

and the prob of selecting a boy in the 3rd round is 3/4

and the prob if selecting a boy in the 4th round is 2/3

So the prob of getting equal girls and boys is 3/6*2/5*3/4*2/3 = 1/10

Can some one pls tell me where I am going wrong???
I make the same mistake sometimes and first try to multiply the individual probabilities and then decide afterwards that something isn't right when I notice the answers choices do not contain my answer, then I go back and rethink the whole problem. This takes a long time!

Thanks in advance.
Join the discussion
Source: — Problem Solving |

I can see your frustration with this Permutations/Combinations problems. You shuold probably check into this resource available on the BTG wiki, exclusively dedicated to this topic:
Here is the link:
https://www.beatthegmat.com/viewtopic.php?t=71

Hope this helps. :!:
A falling tree resounds... but a forest grows in silence...
Join the discussion

wizardofwashington wrote:I can see your frustration with this Permutations/Combinations problems. You shuold probably check into this resource available on the BTG wiki, exclusively dedicated to this topic:
Here is the link:
https://www.beatthegmat.com/viewtopic.php?t=71

Hope this helps. :!:
Thanks, I'll look into it. However, my problem is not necessarily with the theory behind it. I'm an engineering major and aced my combinatorics class in college and I've gone through the Veritas P&C lecture and book. My problem is quickly identifying the solution path.
Join the discussion